### v1

@@ VIII-C2 Ablation Study on the Bias Term $\mathbf{\Lambda}$
(1) Training progression: $\Delta_{\text{cos}}$ tends to increase as training proceeds, which naturally drives the uncertainty scale $s_{u}$ upward in later stages. This behavior aligns with our intuition in Section VIII-C1, that a larger scale is desirable.
(2) Early-stage discrimination: At the early stage of training, the estimated uncertainty $\mathbf{\Sigma}^{s}$ is neither accurate nor sufficiently discriminative across different samples. Incorporating $\Delta_{\text{cos}}$ introduces sample-specific variation into $s_{u}$ for different samples, providing a learning signal that helps the model form more meaningful uncertainty estimates.
@@
To verify the necessity of modeling utterance-level uncertainty, we ablate the covariance term by removing $\mathbf{\Sigma}^{s}$ and retaining only $\mathbf{\Lambda}$ in Equation 16, as shown in Exp. 11. Compared with Exp. 10, this removal leads to a clear performance drop, indicating that estimating utterance-level uncertainty is essential. This result indicates that the gains from global-level uncertainty supervision are not solely due to the dynamic scaling effect, but also arise from more accurate, sample-specific uncertainty estimation provided by $\mathbf{\Sigma}^{s}$.
@@ VIII-D The Effectiveness of Temporal Modeling
In this section, we incorporate a Transformer encoder to model temporal dependencies and compare multi-head attention (MHA) with different head configurations, as reported in Table III. The results show that the 8-head configuration achieves the best performance among all MHA variants. Based on this setting, we further encourage each attention head to capture a distinct temporal context, forming our proposed multi-view self-attention (MVA) mechanism. Experimental results demonstrate that MVA further improves performance, yielding over a 5% relative gain while introducing only minimal additional parameters and computational overhead.
@@ VIII-E $\mathcal{U}^{3}$-xi: Combining All Components
In this section, we integrate all proposed components into a unified model—including the global-level uncertainty supervision ($\mathcal{L}_{\text{AAM}_{u}}$), the multi-view self-attention based uncertainty estimation module, and the uncertainty-aware scoring mechanism—and evaluate the overall performance gain. This combined system, referred to as $\mathcal{U}^{3}$-xi, allows us to examine the cumulative benefit and the interaction among these three complementary designs. Table IV presents the performance improvements obtained by progressively incorporating each component. To more clearly verify the effectiveness of the proposed uncertainty estimation, we use the original ECAPA-TDNN with attentive statistical temporal pooling (ASTP) [16] as the baseline.
@@
(1) Replacing ASTP with a lightweight uncertainty estimation module: The original xi-vector [24] replaces ASTP with a two-layer Gaussian posterior inference module that estimates frame-level uncertainty using two simple linear projections. This architectural change alone yields relative improvements of 6.53% in EER and 8.76% in minDCF, demonstrating the effectiveness of incorporating uncertainty estimation into the speaker embedding framework.
(2) Introducing multi-view self-attention (MVA) for temporal modeling: Building on the previous configuration, MVA-8 incorporates a Transformer encoder with MVA of 8 attention heads to enhance nonlinear modeling capacity and capture richer temporal dependencies. This addition yields further relative performance gains, demonstrating the effectiveness of strengthened temporal uncertainty modeling.
(3) Applying global-level uncertainty supervision: By inserting uncertainty into the scale of softmaxsoftmax function $\mathcal{L}_{\text{AAM}_{u}}$, the model receives explicit supervision to learn more reliable and discriminative uncertainty representations. This global guidance further enhances generalization across both in-domain and cross-domain conditions.
@@
Figure 10: Relationship between utterance-level variance and its duration on the Vox1 test set, including Vox1-O, Vox1-E and Vox1-H.
(4) Incorporating uncertainty scoring: We attempted to learn the scaling factor $\rho$ using the SVL loss, as shown in Table I. However, this approach failed to converge, with NaN values occurring during training, indicating a conflict between $\mathcal{L}_{\text{SVL}}$ and $\mathcal{L}_{\text{AAM}_{u}}$. Consequently, we fix $\rho=1$ as a static scaling factor, which is also intuitively reasonable since the proposed global-level supervision already yields sufficiently discriminative uncertainty estimation. With uncertainty-aware scoring enabled, the final system achieves the largest performance improvement, obtaining relative gains of 21.1% and 15.57% over the baseline in terms of EER and minDCF, respectively.
@@ VIII-F1 Relationship between Uncertainty and Sample Difficulty
We first investigate the relationship between utterance-level uncertainty (i.e., the diagonal covariance $\mathbf{\Sigma}^{s}$) and the cosine gap $\Delta_{\text{cos}}$, which serves as a measure of the classification difficulty of sample $i$, where a larger gap indicates that the sample is easier to classify. Fig. 9 compares this relationship before and after applying our proposed method $\mathcal{U}^{3}$-xi. The Pearson correlation coefficient decreases from -0.166 (xi-vector) to -0.511 ($\mathcal{U}^{3}$-xi), demonstrating that the proposed method strengthens the negative correlation between uncertainty and sample easiness. This trend indicates that $\mathcal{U}^{3}$-xi produces more discriminative and interpretable uncertainty, assigning lower variance to easier samples and higher variance to harder ones. This demonstrates that our global-level uncertainty supervision indeed provides meaningful and effective guidance.
@@ VIII-F2 Relationship between Uncertainty and Sample Duration
Fig. 10 illustrates the relationship between utterance-level uncertainty and utterance duration on the test set, since all training samples are truncated to 2 seconds. Both subfigures show a clear trend that longer utterances tend to have lower estimated variance, which aligns with the intuitive understanding that embeddings extracted from longer speech segments are generally more robust. Interestingly, we observe that this correlation appears weaker for our proposed $\mathcal{U}^{3}$-xi model. Despite this weaker dependency on utterance duration, $\mathcal{U}^{3}$-xi still outperforms the baseline xi-vector model. One possible explanation is that $\mathcal{U}^{3}$-xi may capture aspects of utterance quality beyond length, allowing it to assign lower uncertainty even to short but high-quality utterances.
@@ VIII-F3 The Distribution of Sample-wise Scale Across Training Epochs
Fig. 7 suggests that using a slightly larger scale may benefit optimization. In this section, we further investigate this phenomenon by visualizing how the sample-wise scale evolves across training epochs. Our method predicts a sample-dependent scale, meaning that each utterance is assigned a single scale value during training. For each epoch, we count the number of samples that fall into each scale value and summarize the statistics in Fig. 8. Several observations can be made: 1. In each epoch, some samples receive scales smaller than 32, while others receive scales larger than 32. This indicates that the model adaptively assigns different scales according to sample difficulty. 2. Before the margin starts increasing (epochs $\leq 20$), the mean scale gradually rises. Once the margin begins increasing ($20<\text{epoch}\leq 40$), the mean scale drops significantly. After the margin is fixed (epoch >40>40), the mean scale gradually increases again. This behavior reflects the sensitivity of the scale to changes in the angular margin. At the start of training, the model focuses on easier samples, as smaller scales soften the softmax distribution and cause the gradient to emphasize these samples. As training progresses, the scale gradually increases, shifting the model’s attention toward harder samples. In summary, a dynamic scale plays a crucial role in balancing the contribution of each training utterance. Moreover, the progressively right-shifting scale distribution resembles a curriculum learning strategy [69], guiding the model to focus on easier samples first and gradually adapt to harder ones.
TABLE V: Performance comparison across different speaker encoders. Cells highlighted in gray indicate the application of the proposed methods. ‘LM’ denotes large-margin fine-tuning. ‘AS-Norm’ denotes adaptive S-norm [70]. ‘QMF’ refers to the Quality Measure Function [71]. Results shown in gray correspond to the proposed methods. $\rho=1$ means we apply uncertainty-aware cosine scoring.
Row Speaker 777 indicates that these results are directly inferred using the pretrained WeSpeaker model available at https://github.com/wenet-e2e/wespeaker/blob/master/docs/pretrained.md. means we retrain these models by ourselves. # Param. LM AS- Q In-domain Test Cross-domain Test Avg. $\Delta$
@@
Figure 11: Uncertainty score (Variance) of different utterances changes during training epochs. The margin is linearly increasing during epoch 20 to 40. The y-axis shows the averaged diagonal variance across all embedding dimensions
@@ VIII-F4 Uncertainty Dynamics during Training
From Fig. 11, we observe that the uncertainty values are very small at the early stages of training, and then gradually increase. When the margin starts to increase (epoch 20–40), the uncertainty variance also rises significantly, indicating that the uncertainty estimation module is highly sensitive to changes in the margin. After the margin is fixed, the uncertainty scores gradually decrease as training progresses, which suggests that the network becomes more stable and converges toward a more confident embedding space. This observation implies that, although we aim to model data uncertainty rather than model uncertainty, the estimated uncertainty is still influenced by the optimization status of the model. In addition, the relative uncertainty across different utterances is not consistent throughout training. For example, the uncertainty score of utterance 2 is the highest around epoch 40 but later becomes lower than that of other utterances. This behavior indicates that the uncertainty estimation is not immediately accurate at the beginning of training; rather, it dynamically adjusts and becomes more reliable as training progresses. Overall, no obvious global pattern is observed across all utterances.
@@ VIII-G1 In-domain Test
Considering ECAPA-TDNN512 (Rows 8–10), ResNet34 (Rows 11–13), and ReDimNet-B2 (Rows 14, 18, and 19), we observe that the proposed $\mathcal{U}^{3}$-xi consistently achieves performance improvements over their corresponding baselines. Furthermore, applying uncertainty-aware scoring with $\rho=1$ generally leads to additional gains. The only exceptions are ReDimNet-B2 evaluated on Vox1-E and Vox1-H, where minDCF slightly increases from 0.089 to 0.091 and from 0.153 to 0.157, respectively.
In terms of average performance, the relative improvements are 18.65% for ECAPA-TDNN512, 13.29% for ResNet34, and 9.45% for ReDimNet-B2. These results indicate that while $\mathcal{U}^{3}$-xi remains effective across different speaker encoders, the margin for improvement diminishes as the underlying encoder becomes more advanced.
@@ VIII-G2 Cross-domain Test
The SITW-EVAL set is relatively similar to VoxCeleb in terms of recording conditions; consequently, performance on this dataset follows trends similar to those observed in the in-domain evaluation. Specifically, $\mathcal{U}^{3}$-xi delivers consistent improvements across different speaker encoders, while uncertainty-aware scoring leads to a slight degradation in terms of minDCF.
In contrast, on the more challenging cross-domain CNCeleb dataset, the proposed method improves EER but fails to yield consistent gains in minDCF. Even after applying uncertainty-aware scoring, minDCF further degrades in several cases, with extreme failures observed for Rows 10 and 18, where minDCF spikes to 1.0, clearly indicating invalid result. To further investigate this issue, we plot the Detection Error Tradeoff (DET) curves for Rows 9 and 10, as shown in Fig. 12. It can be observed that the DET curve of Row 10 becomes severely distorted. Compared with the relatively smooth and well-balanced curve of Row 9, the curve of Row 10 exhibits a much steeper and less stable tradeoff between the false-alarm rate (FAR) and the false-reject rate (FRR).
@@
VIII-G3 The effect of LM, AS-Norm and QMF on Uncertainty
We further explore the compatibility of our $\mathcal{U}^{3}$-xi framework with commonly used training-time fine-tuning and backend score processing methods, including LM fine-tuning, AS-Norm, and QMF (Rows 15-17, 20–25 in Table V). Overall, these techniques are compatible with $\mathcal{U}^{3}$-xi and can provide additional performance gains when applied jointly.
Considering the large-margin (LM) setting (Rows 15, 20, and 21), these results illustrate the performance changes after large-margin fine-tuning. Consistent with previous findings, our method improves EER across most test sets, except for Row 20 on CNCeleb. In contrast, while minDCF improves on in-domain and SITW-Eval, it fails on the more challenging cross-domain CNCeleb dataset. Furthermore, after applying uncertainty-aware cosine scoring (Row 21), EER remains improved, whereas minDCF consistently degrades across all test sets.
Considering AS-Norm (Rows 16, 22, and 23), similar trends are observed as with the LM setting: our method performs well on in-domain EER, but shows limited effectiveness on cross-domain data. In addition, uncertainty-aware cosine scoring fails to improve minDCF in both in-domain and cross-domain scenarios.
QMF alleviates these issues to some extent, as seen by comparing Rows 17, 24, and 25. Both Rows 24 and 25 outperform Row 17 on in-domain data in most cases. However, for cross-domain data, the method still fails to achieve improvements.

### v2

@@ VIII-C2 Ablation Study on the Bias Term $\mathbf{\Lambda}$
  1. 1.
    Training progression: $\Delta_{\text{cos}}$ tends to increase as training proceeds, which naturally drives the uncertainty scale $s_{u}$ upward in later stages. This behavior aligns with our intuition in Section VIII-C1, that a larger scale is desirable.
  2. 2.
    Early-stage discrimination: At the early stage of training, the estimated uncertainty $\mathbf{\Sigma}^{\text{s}}$ is neither accurate nor sufficiently discriminative across different samples. Incorporating $\Delta_{\text{cos}}$ introduces sample-specific variation into $s_{u}$ for different samples, providing a learning signal that helps the model form more meaningful uncertainty estimates.
@@
To verify the necessity of modeling utterance-level uncertainty, we ablate the covariance term by removing $\mathbf{\Sigma}^{\text{s}}$ and retaining only $\mathbf{\Lambda}$ in Equation 15, as shown in Exp. 11. Compared with Exp. 10, this removal leads to a clear performance drop, indicating that estimating utterance-level uncertainty is essential. This result indicates that the gains from global-level uncertainty supervision are not solely due to the dynamic scaling effect, but also arise from more accurate, sample-specific uncertainty estimation provided by $\mathbf{\Sigma}^{\text{s}}$.
@@ VIII-D The Effectiveness of Temporal Modeling
In this section, we incorporate a Transformer encoder to model temporal dependencies and compare multi-head attention (MHA) with different head configurations, as reported in Table III. The results show that the 8-head configuration achieves the best performance among all MHA variants. Based on this setting, we further encourage each attention head to capture a distinct temporal context, forming the multi-view self-attention (MVA) mechanism. Experimental results demonstrate that MVA further improves performance, yielding over a 5% relative gain while introducing only minimal additional parameters and computational overhead.
@@ VIII-E $\mathcal{U}^{3}$-xi: Combining All Components
In this section, we integrate all proposed components into a unified model, including the global-level uncertainty supervision ($\mathcal{L}_{\text{UAAM}}$), the MVA uncertainty estimation module, and the uncertainty-aware scoring mechanism. This combined system, referred to as $\mathcal{U}^{3}$-xi, allows us to examine the cumulative benefit and the interaction among these three complementary designs. Table IV presents the performance improvements obtained by progressively incorporating each component. To more clearly verify the effectiveness of the proposed uncertainty estimation, we use the original ECAPA-TDNN with attentive statistical temporal pooling (ASTP) [16] as the baseline.
@@
  1. 1.
    Incorporating uncertainty: The original xi-vector [24] replaces ASTP with a two-layer Gaussian posterior inference module that estimates frame-level uncertainty using two simple linear projections. This architectural change alone yields relative improvements of 6.53% in EER and 8.76% in minDCF, demonstrating the effectiveness of incorporating uncertainty estimation into the speaker embedding framework.
  2. 2.
    Introducing MVA for temporal modeling: Building on the previous configuration, we add a MVA module to enhance nonlinear modeling capacity and capture richer temporal dependencies. This addition yields further relative performance gains, demonstrating the effectiveness of strengthened temporal uncertainty modeling.
  3. 3.
    Applying global-level uncertainty supervision: By inserting uncertainty into the scale of softmaxsoftmax function, $\mathcal{L}_{\text{UAAM}}$, the model receives explicit supervision to learn more reliable uncertainty representations. This global guidance further enhances generalization across both in-domain and cross-domain conditions.
  4. 4.
    Incorporating uncertainty scoring: We initially attempted to learn the scaling factor $\rho$ using the SVL loss in Eq. 10. However, this scheme failed to converge, with NaN values observed, indicating an optimization conflict between $\mathcal{L}_{\text{SVL}}$ and $\mathcal{L}_{\text{UAAM}}$. We therefore fix $\rho=1$ as a static scaling factor. Despite its simplicity, this setting still yields performance gains. This behavior is reasonable, as the proposed global-level uncertainty supervision already enforces discriminative utterance-level uncertainty, making additional learnable scaling unnecessary. With uncertainty-aware scoring enabled, the final system achieves relative improvements of 21.1% and 15.57% over the baseline in terms of EER and minDCF, respectively.
@@ VIII-F Systematic Analysis of Uncertainty in the $\mathcal{U}^{3}$-xi Model
Figure 9: Relationship between utterance-level variance (average variance over embedding dimension) and its duration on the Vox1 test set, including Vox1-O, Vox1-E and Vox1-H.
@@ VIII-F1 Relationship between Uncertainty and Sample Difficulty
We first investigate the relationship between utterance-level uncertainty (i.e., the diagonal covariance $\mathbf{\Sigma}^{\text{s}}$) and the cosine gap $\Delta_{\text{cos}}$, which serves as a measure of the classification difficulty of sample $i$, where a larger gap indicates that the sample is easier to classify. Fig. 8 compares this relationship before and after applying our proposed method $\mathcal{U}^{3}$-xi. The Pearson correlation coefficient decreases from -0.166 (xi-vector) to -0.511 ($\mathcal{U}^{3}$-xi), demonstrating that the proposed method strengthens the negative correlation between uncertainty and sample easiness. This trend indicates that $\mathcal{U}^{3}$-xi produces more discriminative and interpretable uncertainty, assigning lower variance to easier samples and higher variance to harder ones. This demonstrates that our global-level uncertainty supervision indeed provides meaningful and effective guidance.
@@ VIII-F2 Relationship between Uncertainty and Sample Duration
Fig. 9 illustrates the relationship between utterance-level uncertainty and utterance duration on the test set. Both subfigures show a clear trend that longer utterances tend to have lower estimated variance, which aligns with the intuitive understanding that embeddings extracted from longer speech segments are generally more robust. Interestingly, we observe that this correlation appears weaker for our proposed $\mathcal{U}^{3}$-xi model. Despite this weaker dependency on utterance duration, $\mathcal{U}^{3}$-xi still outperforms the baseline xi-vector model. One possible explanation is that $\mathcal{U}^{3}$-xi may capture aspects of utterance quality beyond length, allowing it to assign lower uncertainty even to short but high-quality utterances.
@@ VIII-F3 The Distribution of Sample-wise Scale Across Training Epochs
Fig. 7 suggests that a moderately larger scale can benefit optimization. To further analyze this behavior, we visualize the evolution of sample-wise scales across training epochs. Since each utterance is assigned a single, sample-dependent scale, we summarize the scale distribution at each epoch in Fig. 10. Several observations can be made:
  1. 1.
    Sample-wise perspective: Within each epoch, different samples are assigned different scales. Samples with large scales are typically easy and reliably classified, whereas samples with small scales correspond to hard or unreliable cases. This demonstrates that the model adaptively assigns sample-dependent scales according to the difficulty of each utterance.
  2. 2.
    Epoch-wise perspective: Before the angular margin starts increasing (epochs $\leq 20$), the mean scale gradually rises. When the margin begins to increase ($20<\text{epoch}\leq 40$), the mean scale drops noticeably, indicating the sensitivity of the scale to changes in the margin constraint. After the margin is fixed (epoch >40>40), the mean scale increases again as training proceeds.
    This trend can be explained by the training dynamics. As training progresses, an increasing number of samples become correctly classified, resulting in a higher proportion of easy samples and thus a larger mean scale at each epoch. From a global optimization perspective, a larger average scale sharpens the softmax distribution, increasing gradient sensitivity to misclassified or ambiguous samples.
In summary, the dynamic scale effectively balances the contributions of training utterances. The progressively right-shifting scale distribution further exhibits a curriculum-learning behavior [69], where the model initially emphasizes easier samples and gradually adapts to harder ones.
TABLE V: Performance comparison across different speaker encoders. Cells highlighted in gray indicate the application of the proposed methods. ‘LM’ denotes large-margin fine-tuning. ‘AS-Norm’ denotes adaptive S-norm [70]. ‘QMF’ refers to the Quality Measure Function [71]. Results shown in gray correspond to the proposed methods. $\rho=1$ means uncertainty-aware scoring is applied.
Row Speaker 888 indicates that these results are directly inferred using the pretrained WeSpeaker model available at https://github.com/wenet-e2e/wespeaker/blob/master/docs/pretrained.md. means we retrain these models by ourselves. # Param. LM AS- Q In-domain Test Cross-domain Test Avg. $\Delta$
@@
Figure 11: Uncertainty score (average variance over embedding dimension) of different utterances changes during training epochs. The margin is linearly increasing during epoch 20 to 40. The y-axis shows the averaged diagonal variance across all embedding dimensions
@@ VIII-F4 Uncertainty Dynamics during Training
Fig. 11 illustrates the evolution of estimated uncertainty across training epochs. At early stages, uncertainty values remain low, reflecting the limited discriminative capacity of the model. As training proceeds, uncertainty gradually increases, and rises sharply when the angular margin is introduced (epochs 20–40), indicating that the uncertainty estimation module is highly sensitive to changes in the margin constraint. After the margin is fixed, the uncertainty decreases steadily, suggesting that the embedding space becomes more stable and the model converges toward more confident representations.
Although our objective is to model data (aleatoric) uncertainty rather than model (epistemic) uncertainty, these observations suggest that the estimated data uncertainty is nevertheless influenced by the network’s optimization state. In particular, when the network has not yet converged to a stable solution, accurate estimation of data uncertainty becomes more challenging.
Furthermore, the relative uncertainty among different utterances evolves over time. For instance, an utterance that exhibits high uncertainty at intermediate epochs may later become more confident than others. This behavior suggests that uncertainty estimation is not reliable at early training stages, but gradually adapts and becomes more meaningful as training converges. Overall, no consistent global ranking of uncertainty across utterances is observed throughout training.
@@ VIII-G1 In-domain Test
Considering ECAPA-TDNN512 (Rows 8–10), ResNet34 (Rows 11–13), and ReDimNet-B2 (Rows 14, 18, and 19), we observe that the proposed $\mathcal{U}^{3}$-xi consistently achieves performance improvements over their corresponding baselines. Furthermore, applying uncertainty-aware scoring with $\rho=1$ generally leads to additional gains. The only exceptions are ReDimNet-B2 evaluated on Vox1-E and Vox1-H, where minDCF slightly increases from 0.089 to 0.091 and from 0.153 to 0.157, respectively. In terms of average performance, the relative improvements are 18.64% for ECAPA-TDNN512, 13.29% for ResNet34, and 9.45% for ReDimNet-B2. These results indicate that while $\mathcal{U}^{3}$-xi remains effective across different speaker encoders, the improvement diminishes as the speaker encoder becomes stronger.
@@ VIII-G2 Cross-domain Test
The SITW-EVAL set is relatively similar to VoxCeleb in terms of recording conditions; consequently, its performance trends are consistent with those observed in the in-domain evaluation. Specifically, setting $\rho=0$ yields consistent improvements across different speaker encoders in terms of all metrics, whereas $\rho=1$ results in a slight degradation in terms of minDCF.
In contrast, on the more challenging cross-domain CNCeleb dataset, setting $\rho=0$ fails to achieve stable improvements in either EER or minDCF. Moreover, setting $\rho=1$ consistently improves EER; however, it leads to further degradation in terms of minDCF in several cases, with severe failures observed in Rows 10 and 19, where the minDCF spikes to 1.0, clearly indicating invalid operating points.
To further investigate this issue, we plot the Detection Error Tradeoff (DET) curves for Rows 9 and 10, as shown in Fig. 12. It can be observed that the DET curve of Row 10 becomes severely distorted. Compared with the relatively smooth and well-balanced curve of Row 9, the curve of Row 10 exhibits a much steeper and less stable tradeoff between the false-alarm rate (FAR) and the false-reject rate (FRR).
@@
VIII-G3 Effect of LM, AS-Norm, and QMF on Uncertainty Modeling
We further investigate the compatibility of the proposed $\mathcal{U}^{3}$-xi framework with commonly used training-time fine-tuning and backend score normalization techniques, including large-margin (LM) fine-tuning, AS-Norm, and QMF (Rows 15–17 and 20–25 in Table V). Overall, these techniques remain compatible with $\mathcal{U}^{3}$-xi and can provide additional gains; however, their interaction with uncertainty-aware scoring exhibits distinct behaviors across evaluation settings.
  1. 1.
    Under the LM setting (Rows 15, 20, and 21), fixing $\rho=0$ consistently improves performance over the LM baseline (Row 15) on most evaluation sets, except for the cross-domain CNCeleb dataset. When enabling uncertainty-aware scoring with $\rho=1$, consistent improvements are observed in terms of EER across all test sets, but at the cost of degraded minDCF. We attribute this behavior to the high sensitivity of uncertainty estimation to changes in the angular margin. In particular, the large margin is applied for only a small number of epochs (five epochs in total), which may be insufficient for the uncertainty estimation module to converge to a stable and well-calibrated solution.
  2. 2.
    A similar pattern is observed when AS-Norm is applied (Rows 16, 22, and 23). Setting $\rho=0$ yields consistent improvements over the AS-Norm baseline (Row 16) on most evaluation sets, except for CNCeleb. In contrast, setting $\rho=1$ consistently improves EER while degrading minDCF across all test sets, indicating that AS-Norm does not fully resolve the calibration instability introduced by uncertainty-aware scoring.
  3. 3.
    QMF partially alleviates the above issues, as shown by the comparison among Rows 17, 24, and 25. For in-domain evaluations, both Rows 24 and 25 consistently outperform the QMF baseline (Row 17) on most test sets in terms of either EER or minDCF, even when uncertainty-aware scoring with $\rho=1$ is enabled. However, on cross-domain data, the performance gains remain unstable. Specifically, improvements are only observed on SITW-Eval, with Row 24 improving EER and Row 25 improving minDCF. This suggests that while QMF enhances robustness to uncertainty-aware scoring, domain mismatch remains a fundamental challenge for reliable uncertainty calibration.